OpenStudy (anonymous):

simplify the left side to equal the right side: cos^2x(1+cot^2x)=cot^2x

OpenStudy (anonymous):

help me please

OpenStudy (ashleyisakitty):

I dont know trig!

OpenStudy (anonymous):

ok

OpenStudy (luigi0210):

Change the trig terms

OpenStudy (anonymous):

how?

OpenStudy (luigi0210):

\[\cos^2x(1+\cot^2x)\] \[\cos^2x(\csc^2x)\] \[\cos^2x(\frac{ 1 }{ \sin^2x) }\]

OpenStudy (luigi0210):

multiply it out and guess what you get? :P

OpenStudy (raden):

if u remember the identity of this : 1+cot^2x = csc^2 (x) so, the left side becomes cos^2x * csc^2 (x) = cos^2 x * 1/sin^2 x = cos^2 x / sin^2 x = (cosx/sinx)^2 = cot^2 x QED

OpenStudy (whpalmer4):

If you don't remember your identities (I never do), here it is the basic way: \[\cos^2x(1+\cot^2x) = \cot^2x\]\[\cot x = 1/\tan x = \frac{1}{\frac{\sin x}{\cos x}} = \frac{\cos x}{\sin x}\] \[\cos^2x(1+\frac{\cos^2x}{\sin^2x}) = \frac{\cos^2x}{\sin^2x}\]\[\cos^2x(\frac{\sin^2x}{\sin^2x} + \frac{\cos^2x}{\sin^2x}) = \frac{\cos^2x}{\sin^2x}\]\[\cos^2x(\frac{\sin^2x+\cos^2x}{\sin^2x}) = \frac{\cos^2x}{\sin^2x}\]but \(\sin^2x+\cos^2x=1\) so that becomes \[\cos^2x(\frac{1}{\sin^2x}) = \frac{\cos^2x}{\sin^2x}\]and our work is done.